Solve dy/dx=(1+x)(1+y) please i need assistance to this question... Pls help out

Grade:6

1 Answers

Arun
6 years ago
Dear Student
 
dy /(1+y)  = (1+x) dx
On integrating
ln (1+y) = x + x²/2  + C
 
Regards
Arun (AskIITians forum expert)